# Beyond Standard MoE: Mixture of Latent Experts for Resource-Efficient Language Models

### v1

@@ Abstract
Mixture of Experts (MoE) has emerged as a pivotal architectural paradigm for efficient scaling of Large Language Models (LLMs), operating through selective activation of parameter subsets for each input token. Nevertheless, conventional MoE architectures encounter substantial challenges, including excessive memory utilization and communication overhead during training and inference, primarily attributable to the proliferation of expert modules. In this paper, we introduce Mixture of Latent Experts (MoLE), a novel parameterization methodology that facilitates the mapping of specific experts into a shared latent space. Specifically, all expert operations are systematically decomposed into two principal components: a shared projection into a lower-dimensional latent space, followed by expert-specific transformations with significantly reduced parametric complexity. This factorized approach substantially diminishes parameter count and computational requirements. Beyond the pretraining implementation of the MoLE architecture, we also establish a rigorous mathematical framework for transforming pre-trained MoE models into the MoLE architecture, characterizing the sufficient conditions for optimal factorization and developing a systematic two-phase algorithm for this conversion process. Our comprehensive theoretical analysis demonstrates that MoLE significantly enhances computational efficiency across multiple dimensions while preserving model representational capacity. Empirical evaluations corroborate our theoretical findings, confirming that MoLE achieves performance comparable to standard MoE implementations while substantially reducing resource requirements.
@@ 1 Introduction
Despite their theoretical and empirical advantages, standard MoE architectures (Dai et al., 2024) face significant practical limitations that inhibit broader deployment. These models require substantial memory resources to store parameters across numerous expert modules in Feed-Forward Network (FFN) layers and create communication bottlenecks during distributed training due to all-to-all data transfers. These challenges become increasingly pronounced as models scale to hundreds of experts, potentially limiting their applicability in resource-constrained environments.
Through systematic investigation of parameter utilization in MoE architectures, we discover substantial redundancy within the FFN layers of current MoE models. By analyzing Qwen1.5-MoE-A2.7B (Team, 2024), we reveal that a significant proportion of parameters in its FFN layers can be effectively approximated through lower-dimensional representations without compromising model performance. This empirical observation motivates a fundamental rethinking of expert parameterization in neural architectures.
@@
Figure 1: Architectural comparison between MoE and MoLE in the feed-forward network (FFN) layer. In both diagrams, $N$ denotes the number of experts. MoLE extends the conventional MoE architecture by introducing latent mappings $B_{\text{up}}$, $B_{\text{gate}}$, and $B_{\text{down}}$ that capture shared information across experts. Expert-specific information is encapsulated in the mappings $A_{\text{up}}^{i}$, $A_{\text{down}}^{i}$, and $A_{\text{gate}}^{i}$ for each expert $i$. This latent space decomposition enables MoLE to significantly reduce both memory requirements and computational complexity in FFN layers while maintaining model expressivity.
@@
We introduce Mixture of Latent Experts (MoLE), a novel parameterization paradigm that addresses the core inefficiencies of traditional MoE architectures. Our key insight is that expert modules in standard MoE models contain significant redundancy and operate in unnecessarily high-dimensional spaces. MoLE reformulates each expert operation through a mathematically principled two-phase transformation: (1) a shared projection into a compressed latent space, followed by (2) expert-specific transformations within this lower-dimensional manifold.
@@
Formally, MoLE implements this insight by factorizing each expert’s weight matrix $W^{i}\in\mathbb{R}^{m\times n}$ with $m\ll n$ into the product $A^{i}B$, where $A^{i}\in\mathbb{R}^{m\times m}$ represents expert-specific transformations and $B\in\mathbb{R}^{m\times n}$ represents a shared projection into a latent space of dimension. This factorization yields substantial parameter reduction, particularly in contemporary LLM architectures where the hidden dimension $n$ significantly exceeds the MoE intermediate dimension $m$.
Our contributions are as follows:
  • We propose Mixture of Latent Experts (MoLE), a parameter-efficient architecture that achieves competitive performance with standard MoE models while requiring significantly fewer parameters and reduced computational overhead.
  • We develop a theoretically grounded framework for transforming pre-trained MoE models into the MoLE architecture, including a mathematical characterization of optimal factorization conditions and an efficient two-stage algorithm incorporating low-rank approximation techniques.
@@
  • Through comprehensive empirical evaluation on multiple benchmark datasets, we demonstrate that MoLE preserves or enhances model capabilities across diverse language tasks while substantially improving parameter efficiency, thereby enabling more economical scaling of large language models.
@@ 2 Related Works
Sparse token-choice gating, where the gating function selects top-k experts for each input token, is the most prevalent approach (Fedus et al., 2022; Lepikhin et al., 2020; Zoph et al., 2022). This method is often augmented with auxiliary loss functions to promote balanced expert utilization (Lepikhin et al., 2020; Fedus et al., 2022; Du et al., 2022). Alternative approaches include expert-choice gating, where each expert selects the top-k tokens they will process (Zhou et al., 2022; 2023), and non-trainable gating mechanisms that use predetermined routing strategies (Roller et al., 2021; Costa et al., 2022; Gururangan et al., 2021).
A promising recent development in MoE research is the integration with parameter-efficient fine-tuning (PEFT) techniques, creating Mixture of Parameter-Efficient Experts (MoPEs) (Zhang et al., 2021; Wu et al., 2022; Ye et al., 2023). These approaches combine the task versatility of MoE with the resource efficiency of PEFT, positioning them as a significant advancement in efficient multi-task learning.
3 MoLE: A Novel Approach to Expert-based Model Architecture
In this paper, we introduce the Mixture of Latent Experts (MoLE), a novel framework that maps experts into latent space to address several limitations of traditional Mixture of Experts (MoE) architectures. For clarity and focus, we exclude shared-experts from our analysis throughout this paper.
3.1 Background: Finer-Grained MoE Architecture
Finer-grained MoE architectures for Feed-Forward Networks (FFNs) employ $N$ distinct experts (Dai et al., 2024). For each expert $E_{i}(x)$ where $i\in\{1,2,\ldots,N\}$, the computation is defined as:
$E_{i}(x):=W^{i}_{\text{down}}\left(W^{i}_{\text{up}}(x)\odot\text{Act}\left(W_{\text{gate}}(x)\right)\right),$ (3.1)
@@ 3.2 Mixture of Latent Experts: Concept and Design
To address these limitations, we propose the Mixture of Latent Experts (MoLE) framework, which fundamentally reimagines the structure of FFN layers in expert-based systems.
Our approach is informed by a careful analysis of the expert computation in Equation equation 3.1, which comprises three distinct operations:
@@
A critical insight from this analysis is that the core functionality of experts primarily stems from the non-linear transformation in the second step. The projection operations in the first and third steps primarily serve to reduce computational complexity, potentially at the cost of limiting the expert’s specialized knowledge capacity.
Drawing inspiration from both Multi-Head Latent Attention (MLA) (Liu et al., 2024), which introduces a “latent space” for KV caches in attention layers, and Grouped-Query Attention (GQA) (Ainslie et al., 2023), which leverages group-based processing, we propose the Mixture of Latent Experts (MoLE). This architecture fundamentally reconsiders how inputs are mapped to a lower-dimensional latent space where expert computation occurs more efficiently.
@@ 3.3 Formulation of Mixture of Latent Experts
$W_{\text{up}}^{i}=A_{\text{up}}^{i}B_{\text{up}}.$ (3.3)
In this formulation, $B_{\text{up}}\in\mathbb{R}^{m\times n}$ functions as a unified projection operator shared across experts, mapping inputs from the high-dimensional space $\mathbb{R}^{n}$ to a lower-dimensional latent space $\mathbb{R}^{m}$, where typically $m\ll n$. Conversely, $A_{\text{up}}^{i}\in\mathbb{R}^{m\times m}$ represents an expert-specific linear transformation within this latent space, encapsulating the specialized function of expert $i$. Following the terminology established in Multi-Head Latent Attention (MLA), we designate $B_{\text{up}}$ as the latent mapping for the “up operator”.
This factorization principle extends systematically to the “gate operator” as well. On the other side, for the “down operator” $W_{\text{down}}^{i}$, which maps from a lower-dimensional to a higher-dimensional space, the decomposition necessarily assumes a reverse form: $W_{\text{down}}^{i}=B_{\text{down}}A_{\text{down}}^{i}$, where $A_{\text{down}}^{i}\in\mathbb{R}^{m\times m}$ and $B_{\text{down}}\in\mathbb{R}^{n\times m}$.
To optimize the trade-off between model expressivity and parameter efficiency, we introduce a structured grouping mechanism where each subset of $k$ experts shares the same latent mapping matrices $B_{\text{up}}$ and $B_{\text{down}}$. This design establishes a configurable spectrum of architectural possibilities: when $k=1$, each expert maintains its independent latent space, and MoLE becomes functionally equivalent to the standard MoE architecture. Conversely, as $k$ increases, the model achieves progressively higher parameter efficiency at a measured trade-off with expert specialization. This parameterization allows for systematic exploration of the efficiency-performance frontier in mixture-of-experts architectures.
3.4 Formal Definition of MoLE
@@
$E_{i}(x)=B_{\text{down}}^{\lfloor i/k\rfloor}A_{\text{down}}^{i}\left(A_{\text{up}}^{i}B_{\text{up}}^{\lfloor i/k\rfloor}(x)\odot\text{Act}\left(A_{\text{gate}}^{i}B_{\text{gate}}^{\lfloor i/k\rfloor}(x)\right)\right).$ (3.4)
Consequently, the output of the FFN layer employing the MoLE architecture is computed as:
$y=x+\sum_{i=1}^{N}g_{i}(x)E_{i}(x).$ (3.5)
This formulation effectively disentangles expert-specific computations from the shared dimensionality reduction operations, enabling significant parameter efficiency while preserving model expressivity. The visual comparison between MoE and MoLE architectures is shown in Figure 1.
3.5 Efficiency Benefits of MoLE
The MoLE architecture offers significant efficiency advantages over standard fine-grained MoE models, spanning multiple computational dimensions from memory usage to communication overhead. To quantify these benefits, we provide a comparative analysis between MoE and MoLE for a single FFN layer, assuming identical configurations with hidden dimension $n$, MoE intermediate dimension $m$, and number of experts $N$.
@@ 3.5 Efficiency Benefits of MoLE
MoLE (Ours) $3Nm^{2}+3\lfloor\frac{N}{k}\rfloor mn$ $\left(3m^{2}+2m\right)N+3\lfloor\frac{N}{k}\rfloor mn$
Table 1: Efficiency comparison between standard MoE and our proposed MoLE architectures for a single FFN layer. Parameter count represents the total number of trainable parameters in expert modules, while FLOPs measures the computational cost during a forward pass with top-1 routing. Here, $n$ represents the hidden dimension, $m$ the MoE intermediate dimension, $N$ the number of experts, and $k$ the number of latent experts per group.
@@
Parameter Efficiency As shown in Table 1, MoLE substantially reduces the parameter count compared to standard MoE, particularly when $m\ll n$, which is the typical case in modern LLMs. For instance, in DeepSeek-V3, $n=7168$ while $m=2048$. The parameter reduction stems from our latent parameterization, where expert-specific transformations operate in the lower-dimensional latent space ($m\times m$) rather than directly on the high-dimensional hidden space ($m\times n$).
@@
Computational Efficiency Beyond parameter savings, MoLE also reduces the computational cost measured in FLOPs. The efficiency gain becomes particularly pronounced when the number of experts $N$ is large and $k$ is small (meaning fewer latent projection matrices are used). This computational advantage translates to faster inference and training times, especially on hardware where memory bandwidth is a bottleneck.
@@
Communication Overhead Reduction A critical but often overlooked benefit of MoLE is the reduction in all-to-all communication costs during distributed training and inference. In standard MoE models, the full expert parameters ($3Nmn$ in total) must be synchronized across devices. In contrast, MoLE requires synchronization of significantly fewer parameters, reducing network bandwidth requirements and improving scalability for distributed deployments.
@@
Memory Access Patterns MoLE also offers improved cache efficiency during computation. The smaller matrices used in latent transformations ($A^{i}\in\mathbb{R}^{m\times m}$) exhibit better locality of reference compared to the larger matrices in standard MoE ($W^{i}\in\mathbb{R}^{m\times n}$), potentially leading to higher utilization of fast cache memory and reduced main memory bandwidth demands.
4 Theoretical Analysis of MoE to MoLE Transformation

### v2

# MoLAE: Mixture of Latent Experts for Parameter-Efficient Language Models

@@ Abstract
Mixture of Experts (MoE) has become a key architectural paradigm for efficiently scaling Large Language Models (LLMs) by selectively activating a subset of parameters for each input token. However, standard MoE architectures face significant challenges, including high memory consumption and communication overhead during distributed training. In this paper, we introduce Mixture of Latent Experts (MoLAE), a novel parameterization that addresses these limitations by reformulating expert operations through a shared projection into a lower-dimensional latent space, followed by expert-specific transformations. This factorized approach substantially reduces parameter count and computational requirements, particularly in existing LLMs where hidden dimensions significantly exceed MoE intermediate dimensions. We provide a rigorous mathematical framework for transforming pre-trained MoE models into MoLAE architecture, characterizing conditions for optimal factorization, and developing a systematic two-step algorithm for this conversion. Our comprehensive theoretical analysis demonstrates that MoLAE significantly improves efficiency across multiple dimensions while preserving model capabilities. Experimental results confirm that MoLAE achieves comparable performance to standard MoE with substantially reduced resource requirements.
@@ 1 Introduction
Despite their theoretical and empirical advantages, standard MoE architectures (Dai et al., 2024) face significant practical limitations that inhibit broader deployment. These models require substantial memory resources to store parameters across numerous expert modules in Feed-Forward Network (FFN) layers and create communication bottlenecks during distributed training due to all-to-all data transfers. These challenges become increasingly pronounced as models scale to hundreds of experts, potentially limiting their applicability in resource-constrained environments. Through systematic investigation of parameter utilization in MoE architectures, we discover substantial redundancy within the FFN layers of current MoE models. By analyzing Qwen1.5-MoE-A2.7B (Team, 2024), we reveal that a significant proportion of parameters in its FFN layers can be effectively approximated through lower-dimensional representations without compromising model performance. This empirical observation motivates a fundamental rethinking of expert parameterization in neural architectures.
@@
Figure 1: Architectural comparison between MoE and MoLAE in the FFN layer. In both diagrams, $N_{r}$ denotes the number of routed experts. MoLAE extends the conventional MoE architecture by introducing latent mappings $B_{\text{up}}$, $B_{\text{gate}}$, and $B_{\text{down}}$ that capture shared information across experts. Expert-specific information is encapsulated in the mappings $A_{\text{up}}^{i}$, $A_{\text{down}}^{i}$, and $A_{\text{gate}}^{i}$ for each expert $i$.
@@
In this work, we introduce Mixture of LAtent Experts (MoLAE), a novel parameterization paradigm that addresses the core inefficiencies of traditional MoE architectures. Our key insight is that expert modules in standard MoE models contain significant redundancy and operate in unnecessarily high-dimensional spaces. MoLAE reformulates each expert operation through a mathematically principled two-phase transformation: (1) a shared projection into a compressed latent space, followed by (2) expert-specific transformations within this lower-dimensional manifold.
@@
Formally, MoLAE implements this insight by factorizing each expert’s weight matrix $W^{i}\in\mathbb{R}^{m\times n}$ into the product $A^{i}B$, where $A^{i}\in\mathbb{R}^{m\times m}$ represents expert-specific transformations and $B\in\mathbb{R}^{m\times n}$ represents a shared projection into a latent space of dimension. This factorization yields substantial parameter reduction, particularly in contemporary LLM architectures where the hidden dimension $n$ significantly exceeds the MoE intermediate dimension $m$.
@@
Our contributions are as follows: 1) We propose MoLAE, a parameter-efficient architecture that achieves competitive performance with standard MoE models while requiring significantly fewer parameters and reduced computational overhead. 2) We develop a theoretically grounded framework for transforming pre-trained MoE models into the MoLAE architecture, including a mathematical characterization of optimal factorization conditions and an efficient two-stage algorithm incorporating low-rank approximation techniques. 3) Through comprehensive empirical evaluation on multiple benchmark datasets, we demonstrate that MoLAE preserves or enhances model capabilities across diverse language tasks while substantially improving parameter efficiency, thereby enabling more economical scaling of large language models.
@@ 2 Related Works
Sparse token-choice gating, where the gating function selects top-k experts for each input token, is the most prevalent approach (Fedus et al., 2022; Lepikhin et al., 2020; Zoph et al., 2022). This method is often augmented with auxiliary loss functions to promote balanced expert utilization (Lepikhin et al., 2020; Fedus et al., 2022; Du et al., 2022). Alternative approaches include expert-choice gating, where each expert selects the top-k tokens they will process (Zhou et al., 2022, 2023), and non-trainable gating mechanisms that use predetermined routing strategies (Roller et al., 2021; Costa et al., 2022; Gururangan et al., 2021).
A promising recent development in MoE is the integration with parameter-efficient fine-tuning (PEFT) techniques, creating Mixture of Parameter-Efficient Experts (MoPEs) (Zhang et al., 2021; Wu et al., 2022; Ye et al., 2023). These approaches combine the task versatility of MoE with the resource efficiency of PEFT, positioning them as a significant advancement in efficient multi-task learning.
3 Redundancy in Standard MoE models
3.1 Background: Standard MoE Architecture
Finer-grained MoE architectures for FFNs employ $N$ distinct experts (Dai et al., 2024). For each expert $E_{i}(x)$ where $i\in\{1,2,\ldots,N\}$, the computation is defined as:
$E_{i}(x):=W^{i}_{\text{down}}\left(W^{i}_{\text{up}}(x)\odot\textsc{Act}\left(W_{\text{gate}}(x)\right)\right),$ (1)
@@ 3.1 Background: Standard MoE Architecture
3.2 Parameter Redundancy in MoE Models
In this section, we provide empirical evidence for significant parameter redundancy within FFN layers, substantiating the theoretical framework presented in Section 5. We conduct analysis using the Qwen1.5-MoE-A2.7B (Team, 2024) model, a popular MoE model comprising 14.3B parameters while activating only 2.7B parameters during inference, with 60 distinct experts.
For our analysis, we define a ratio-$r$ low-rank approximation $\tilde{W}$ of any matrix $W$ as a matrix whose rank equals $r\times\text{rank}(W)$, where $0<r\leq 1$. In accordance with the Eckart-Young-Mirsky theorem (Schmidt, 1989), these approximations are computed via Singular Value Decomposition (SVD), retaining only the largest $r\times\text{rank}(W)$ singular values and their corresponding singular vectors. To rigorously assess model capabilities under low-rank constraints, we evaluate performance on three benchmark tasks: MMLU (Hendrycks et al., 2021), GSM8K (Cobbe et al., 2021), and Wikitext-2 (Merity et al., 2016). All experiments were conducted using the lm-evaluation-harness evaluation framework (Gao et al., 2024). Table 1 presents the comparative results.
Table 1: Performance comparison of MoE models with varying low-rank approximation ratios across multiple benchmarks.
$1.0$ (Original) 60.1 61.0 9.49
$0.8$ 61.2 60.3 9.65
$0.6$ 60.1 59.4 9.85
Table 1 illustrates the relationship between rank reduction and model performance. The baseline case ($r=1.0$) represents the original, unmodified model with full-rank weight matrices. Notably, when reducing the rank of FFN operators by 20% ($r=0.8$), we observe no significant performance degradation. In fact, the reduced-rank model demonstrates superior performance on the GSM8K benchmark, achieving a 1.1 percentage point improvement over the original model, while maintaining comparable performance on MMLU accuracy and Wikitext-2 PPL. These empirical findings provide compelling evidence that, despite the mathematical full-rank property of FFN weight matrices, a substantial proportion of parameters contain redundant information that can be effectively approximated through lower-dimensional representations. This parameter redundancy phenomenon forms the empirical foundation for our theoretical analysis in Section 5.
4 MoLAE: Mixture of Latent Experts
In this section, we introduce the Mixture of Latent Experts (MoLAE), a novel framework that maps experts into latent space to address several limitations of traditional MoE models. For clarity and focus, we exclude shared-experts from our analysis throughout this paper.
@@ 4.1 Mixture of Latent Experts: Concept and Design
To address these limitations, we propose MoLAE framework, which fundamentally redefines the structure of FFN layers in expert-based systems. Our approach is informed by a careful analysis of the expert computation in Equation (1), which comprises three distinct operations:
@@
A critical insight is that the core functionality of experts primarily stems from the non-linear transformation in the second step. The projection operations in the first and third steps primarily serve to reduce computational complexity, potentially at the cost of limiting the expert’s domain capacity.
Drawing inspiration from both Multi-Head Latent Attention (MLA) (Liu et al., 2024), which introduces a “latent space” for KV caches in attention layers, and Grouped-Query Attention (GQA) (Ainslie et al., 2023), which leverages group-based processing, we propose the MoLAE, which operates on experts, i.e. specific FFN layers in standard MoE models. This architecture fundamentally reconsiders how inputs are projected into a lower-dimensional latent space, enabling more efficient computation within the experts.
@@ 4.2 Formulation of Mixture of Latent Experts
$W_{\text{up}}^{i}=A_{\text{up}}^{i}B_{\text{up}}.$ (3)
In this formulation, $B_{\text{up}}\in\mathbb{R}^{m\times n}$ functions as a unified projection operator shared across experts, mapping inputs from the high-dimensional space $\mathbb{R}^{n}$ to a lower-dimensional latent space $\mathbb{R}^{m}$, where typically $m\ll n$. Conversely, $A_{\text{up}}^{i}\in\mathbb{R}^{m\times m}$ represents an expert-specific linear transformation within this latent space, encapsulating the specialized function of expert $i$. Following the terminology established in MLA, we designate $B_{\text{up}}$ as the latent mapping for the “up operator”. This factorization principle extends systematically to the “gate operator” as well. On the other side, for the “down operator” $W_{\text{down}}^{i}$, which maps from a lower-dimensional to a higher-dimensional space, the decomposition necessarily assumes a reverse form: $W_{\text{down}}^{i}=B_{\text{down}}A_{\text{down}}^{i}$, where $A_{\text{down}}^{i}\in\mathbb{R}^{m\times m}$ and $B_{\text{down}}\in\mathbb{R}^{n\times m}$.
To optimize the trade-off between model expressivity and parameter efficiency, we introduce a structured grouping mechanism where each subset of $k$ experts shares the same latent mapping matrices $B_{\text{up}}$ and $B_{\text{down}}$. This design establishes a configurable spectrum of architectural possibilities: when $k=1$, each expert maintains its independent latent space, and MoLAE becomes functionally equivalent to the standard MoE architecture. Conversely, as $k$ increases, the model achieves progressively higher parameter efficiency at a measured trade-off with expert specialization. This parameterization allows for systematic exploration of the efficiency-performance frontier in mixture-of-experts architectures.
@@
$E_{i}(x)=B_{\text{down}}^{\lfloor i/k\rfloor}A_{\text{down}}^{i}\left(A_{\text{up}}^{i}B_{\text{up}}^{\lfloor i/k\rfloor}(x)\odot\textsc{Act}\left(A_{\text{gate}}^{i}B_{\text{gate}}^{\lfloor i/k\rfloor}(x)\right)\right),$ (4)
where $k$ is the group size of experts. Consequently, the output of the FFN layer employing the MoLAE architecture is computed as:
$y=x+\sum_{i=1}^{N}g_{i}(x)E_{i}(x).$ (5)
This formulation effectively disentangles expert-specific computations from the shared dimensionality reduction operations, enabling significant parameter efficiency while preserving model expressivity. The visual comparison between MoE and MoLAE architectures is shown in Figure 1.
4.3 Efficiency Benefits of MoLAE
The MoLAE architecture offers significant efficiency advantages over standard MoE models, spanning multiple computational dimensions from memory usage to communication overhead. To quantify these benefits, we provide a comparative analysis between MoE and MoLAE for a single FFN layer, assuming identical configurations with hidden dimension $n$, MoE intermediate dimension $m$, and number of experts $N$.
Table 2: Efficiency comparison between standard MoE and our proposed MoLAE architectures for a single FFN layer.
@@ 4.3 Efficiency Benefits of MoLAE
MoLAE (Ours) $3Nm^{2}+3\lfloor\frac{N}{k}\rfloor mn$ $\left(3m^{2}+2m\right)N+3\lfloor\frac{N}{k}\rfloor mn$
@@
Parameter Efficiency As shown in Table 2, MoLAE substantially reduces the parameter count compared to standard MoE, particularly when $m\ll n$, which is the typical case in modern LLMs. For instance, in DeepSeek-V3, $n=7168$ while $m=2048$. The parameter reduction stems from our latent parameterization, where expert-specific transformations operate in the lower-dimensional latent space ($m\times m$) rather than directly on the high-dimensional hidden space ($m\times n$).
@@
Computational Efficiency Beyond parameter savings, MoLAE also reduces the computational cost measured in FLOPs. The efficiency gain becomes particularly pronounced when the number of experts $N$ is large and $k$ is small (meaning fewer latent projection matrices are used). This computational advantage translates to faster inference and training times, especially on hardware where memory bandwidth is a bottleneck.
@@
Communication Overhead Reduction A critical but often overlooked benefit of MoLAE is the reduction in all-to-all communication costs during distributed training and inference. In standard MoE models, the full expert parameters ($3Nmn$ in total) must be synchronized across devices. In contrast, MoLAE requires synchronization of significantly fewer parameters, reducing network bandwidth requirements and improving scalability for distributed deployments.
@@
Memory Access Patterns MoLAE also offers improved cache efficiency during computation. The smaller matrices used in latent transformations ($A^{i}\in\mathbb{R}^{m\times m}$) exhibit better locality of reference compared to the larger matrices in standard MoE ($W^{i}\in\mathbb{R}^{m\times n}$), potentially leading to higher utilization of fast cache memory and reduced main memory bandwidth demands.
5 Transformation from MoE to MoLAE